# Simultaneous elections make single-party sweeps more likely

### v1

@@ Abstract
In a country with multiple elections, it may be expedient to hold some or all of them on a common polling date. Our main result, Theorem C, is that under certain assumptions, an increase in the simultaneity of polling increases the likelihood of a sweep, i.e. the likelihood that a single party wins all the elections. We discuss the applicability of our result to the two most common real world electoral systems, namely first-past-the-post (most voters, including US and India) and party list proportional representation (most countries).
We deduce Theorem C from a certain inequality proved in Theorem D, which is of independent interest. In particular, we connect our inequality to the Harris correlation inequality, which is a multivariate generalization of the Chebyshev sum inequality, and plays an important role in statistical mechanics and graph theory. More precisely, we show that Theorem D also implies Theorem F, which extends the domain of the Harris inequality to a larger class of functions.
@@
Keywords: Simultaneous elections, aligned functions, Harris inequality.
@@ 1 Introduction
Our main theme is that, under certain assumptions, a more simultaneous polling schedule makes it more likely that one party wins all the elections. We make this precise in three scenarios with increasing degree of generality.
@@ 1.1 Main results
Country A has two political parties and two national elections, Presidential and Parliamentary; in each election the candidate with the most votes wins. The election commission is considering whether to hold the elections simultaneously or on two separate polling dates. Each voter $h$ has a fixed probability $p_{h}$ of voting on any poll date [^1] This assumption is very much in the spirit of Banzhaf [4] and Penrose [30]., and the population is polarized in the sense that each voter prefers the same party for both elections. We define a sweep to be the outcome in which the same party wins both the presidency and a parliamentary majority.
@@
The simultaneous election has a higher222throughout, “higher” is to be understood in the weak sense of “greater than or equal to.” likelihood of a sweep.
Country B has several political parties and several elections with varying notions of what constitutes a “win”. As before, each voter has a fixed probability of voting on any poll date and always votes for the same party. Moreover, the following two conditions hold for each election:
(a) for any voter turnout, at most one party can win;
(b) for any voter turnout, an additional vote for one party cannot turn its win into a non-win, nor turn some other party’s non-win into a win.
The election commission is exploring various polling schedules. We will say that schedule $\pi$ is coarser than schedule $\pi^{\prime}$ if any two elections that are simultaneous (have the same poll date) in $\pi^{\prime}$ are also simultaneous in $\pi$. We (re)-define a sweep to be the scenario in which one party wins all the elections.
@@
A coarser polling schedule has a higher likelihood of a sweep.
Country C has a setup entirely similar to country B. However the election commission is considering staggered schedules in which each election might be spread out over several days, with different poll dates for different regions. We will say that staggered schedule $\Pi$ is coarser than $\Pi^{\prime}$ if any two elections that are simultaneous for a voter $h$ in $\Pi^{\prime}$ are also simultaneous for $h$ in $\Pi$.
@@
A coarser staggered schedule has a higher likelihood of a sweep.
In a country with many political parties, it often happens that several parties come together in an alliance or coalition in order to form a government. Here one can distinguish two kinds of alliances – a pre-poll alliance which involves seat sharing, and a post-poll alliance in which a group of parties comes together after having fought the elections separately, perhaps compteing against each other for more seat. In this case, one may ask whether Theorem C continues to hold if we treat each coalition as single party? In section 5.6 we show that is indeed the case for many real -world electoral systems.
@@ 1.2 Structure of the paper
In section 5 we discuss the applicability of our results to various electoral systems, especially the US and Indian electoral systems, and contrast our results to other important themes in the literature such as social choice theory [3, 14].
As for the proofs, we note first that Theorem C implies Theorems A and B. In section 2 we reduce the proof of Theorem C to a mathematical result, Theorem D, which we then prove in section 3. Finally, in section 4, we discuss some further consequences of Theorem D, which might be of independent interest. This includes Theorem F which generalizes [^3] Theorem D generalizes the Harris inequality to correlations involving a larger class of functions. This is in contrast to the well-known FKG inequality [15] which extends the Harris inequality to a larger class of probability measures. One of the authors of the present paper has proposed a further generalization of the FKG inequality to several functions and obtained some partial results in this direction. However the general problem remains open, and seems quite hard; see [22, 33, 34, 35, 36] the Harris inequality – a well-known correlation inequality with applications in many different areas such as statistical mechanics, percolation theory, graph theory, and combinatorics [2, 17].
2 Proof of Theorem C
Let $H=\{1,\ldots,m\}$ and let $L=\{1,\ldots,n\}$ be the sets of voters and elections. The voter turnout, $T\subset H\times L,$ is the set of pairs $(h,l)$ such that voter $h$ has voted in election $l$. This turnout can be equivalently described by the tuple $(L_{1},\ldots,L_{m})$ where $L_{h}\subset L$ is the set of elections in which voter $h$ has voted, or by the tuple $(H_{1},\ldots,H_{n})$ where $H_{l}\subset H$ is the set of voters who have voted in election $l$.
Indeed, if $\mathcal{H},\;\mathcal{L}$ and $\mathcal{T}$ are the power sets (sets of all subsets) of $H,\;L$ and $H\times L$, then we have natural bijections $\lambda:\mathcal{L}^{m}\rightarrow\mathcal{T}$ and $\theta:\mathcal{H}^{n}\rightarrow\mathcal{T}$ given by
$\lambda(L_{1},\ldots,L_{m})=\bigcup\nolimits_{h}\{(h,l):l\in L_{h}\},\quad\theta(H_{1},\ldots,H_{n})=\bigcup\nolimits_{l}\{(h,l):h\in H_{l}\}.$ (1)
A staggered schedule determines a sequence $\Pi=(\pi_{1},\ldots,\pi_{m})$, where $\pi_{h}=\{M^{h}_{1},M^{h}_{2},\ldots\}$ is a partition of $L$ into subsets grouped by common poll dates for $h$. The probabilities $\mathbf{p}=(p_{1},\ldots,p_{m})$ induce a probability measure on $\mathcal{T}$ as follows.
@@ 2 Proof of Theorem C
We write $\mu_{\Pi,\mathbf{p}}(T)$ for the probability of obtaining $T$ by the following random procedure [^4] It is not too difficult to give an explicit product formula for $\mu_{\Pi,\mathbf{p}}(T)$ in terms of $\lambda^{-1}(T)=(L_{1},\ldots,L_{m})$, but we choose to skip this since it will not play any role in the proof.: For each $h\in H$ and for each $M^{h}_{i}$ in $\pi_{h}$, toss a coin with probability $p_{h}$ of heads, and if heads then include the set $\{(h,l):l\in M^{h}_{i}\}$ in $T$.
We define the “sweep” function for party $t$ on $\mathcal{T}$ as follows:
$X^{t}:T\xrightarrow{\theta^{-1}}(H_{1},\ldots,H_{n})\mapsto\chi_{1}^{t}(H_{1})\cdots\chi_{n}^{t}(H_{n}).$ (2)
where $\chi_{l}^{t}(H^{\prime})=1$ if party $t$ wins election $l$ when voters from $H^{\prime}$ cast their votes, and $\chi_{l}^{t}(H^{\prime})=0$ otherwise. Then turnout $T$ yields a sweep for party $t$ if and only if we have $X^{t}(T)=1$, and the probability of a sweep by party $t$ is the expectation
$\sum\nolimits_{T}X^{t}(T)\mu_{\Pi,\mathbf{p}}(T).$
We say a real function $f$ on $\mathcal{H}$ is increasing (resp. decreasing) at $h\in H$ if for all $S\subset H$ one has $f(S\cup\{h\})\geq f(S)$ (resp. $f(S\cup\{h\})\leq f(S)$). By our assumptions the functions $\chi^{t}_{1},\ldots,\chi^{t}_{n}$ are all increasing at $h$ if voter $h$ prefers party $t$, otherwise they are all decreasing at $h$. This motivates the following definition.
@@
We say that the functions $f_{1},\ldots,f_{n}$ on $\mathcal{H}$ are aligned if for each $h\in H$, either they are all increasing at $h$ or all decreasing at $h$.
In the next section we will prove the following result.
@@
Suppose $f_{1},\ldots,f_{n}$ are aligned and nonnegative on $\mathcal{H}$, and define
$F:T\xrightarrow{\theta^{-1}}(H_{1},\ldots,H_{n})\mapsto f_{1}(H_{1})\cdots f_{n}(H_{n}),\quad E(\Pi,\mathbf{p})=\sum\nolimits_{T}F(T)\mu_{\Pi,\mathbf{p}}(T).$
If $\;\Pi^{\prime}$ is coarser than $\Pi$ then we have $E(\Pi^{\prime},\mathbf{p})\geq E(\Pi,\mathbf{p})$.
Proof. (of Theorem C.) By Theorem D and the preceding discussion it follows that if we switch from $\Pi$ to $\Pi^{\prime}$, then this increases the probability of a sweep by party $t$ for every $t$. This implies Theorem C.
3 Proof of Theorem D
Proof. (of Theorem D.) It suffices to prove $E(\Pi^{\prime},\mathbf{p})\geq E(\Pi,\mathbf{p})$ for the special case where
  1. 1.
@@ 3 Proof of Theorem D
where $p=p_{1}$. Now by an easy calculation we get
@@
4 Further consequences of Theorem D
For $n=2$, the set $L=\{1,2\}$ has only two partitions, namely $\{L\}$ and $\{\{1\},\{2\}\}$. In this setting, if $S$ is a subset of $H$ then we can define a partition sequence $\Pi_{S}=(\pi_{h})$ of $L$ as follows: $\pi_{h}=\{L\}$ if $h\in S$, and $\pi_{h}=\{\{1\},\{2\}\}$ if $h\notin S$.
The power set $\mathcal{T}$ of $H\times L$ can be identified with $\mathcal{H}\times\mathcal{H}$, and we write $\mu_{S}$ for the measure $\mu_{\Pi_{S},\mathbf{p}}$. Equivalently $\mu_{S}(H_{1},H_{2})$ is the probability of generating the pair $(H_{1},H_{2})$ by the following random procedure: For each $h\in S$ toss a coin with probability $p_{h}$ of heads, and if heads include $h$ in both $H_{1}$ and $H_{2}$, for $h\notin S$ toss the coin twice, once for inclusion in $H_{1}$ and once for $H_{2}$.
Definition 3
For functions $f_{1}$ and $f_{2}$ on $\mathcal{H}$ we define their convolution $f_{1}\ast f_{2}$ as follows: $f_{1}\ast f_{2}(S)=\sum_{\mathcal{H}\times\mathcal{H}}f_{1}(H_{1})f_{2}(H_{2})\mu_{S}(H_{1},H_{2})$.
The following theorem generalizes a result from [9] for increasing functions.
@@ 4 Further consequences of Theorem D
If $f_{1},f_{2}$ are aligned and nonnegative then $f_{1}\ast f_{2}$ is increasing.
Proof. In the notation of Theorem D for $n=2$ we have $f_{1}\ast f_{2}(S)=E(\Pi_{S},\mathbf{p})$. Evidently $\Pi_{S^{\prime}}$ is coarser than $\Pi_{S}$ if and only if $S^{\prime}$ contains $S$. Thus $f_{1}\ast f_{2}$ is an increasing function.
Let $\mu$ be the product measure on $\mathcal{H}$ defined by
$\mu(S)=\prod_{h\in S}(p_{h})\prod_{h\notin S}(1-p_{h})$
and let $E(f,\mu)=\sum_{S}f(S)\mu(S)$ be the corresponding expected value. Then the following result generalizes the Harris inequality [17] for increasing functions.
Theorem F
If $f_{1},f_{2}$ are aligned and nonnegative then we have
$E(f_{1}f_{2},\mu)\geq E(f_{1},\mu)E(f_{2},\mu).$
Proof. A straightforward calculation shows that we have $f_{1}\ast f_{2}(H)=E(f_{1}f_{2},\mu)$ and $f_{1}\ast f_{2}(\emptyset)=E(f_{1},\mu)E(f_{1},\mu)$. Thus the result follows from Theorem E.

### v2

# Simultaneous elections in a polarized society make single-party sweeps more likely

@@ Abstract
In a country with many elections, it may prove economically expedient to hold multiple elections simultaneously on a common polling date. We show that in a polarized society, in which each voter has a preferred party, an increase in the simultaneity of polling will increase the likelihood of a single-party sweep, namely, it will become more likely that a single party wins all the elections.
In fact we show that the sweep probability goes up for every party. Thus the phenomenon we describe is independent of the “coattail” or “down-ballot” effect of a popular leader. It is a systemic and persistent macroscopic political change, effected by a combination of political polarization and simultaneity of polling.
Our result holds under fairly general conditions and is applicable to many common real-world electoral systems, including first-past-the-post (most voters) and party list proportional representation (most countries). In the course of our proof, we obtain a generalization of the well-known Harris correlation inequality.
@@
Keywords: Simultaneous elections, single-party sweeps, Harris inequality.
@@ 1 Introduction
Most major democracies around the world conduct multiple elections, allowing citizens to engage in the democratic process at regular intervals across various levels of government. Americans participate in elections every two years to choose a member of the House of Representatives, every four years to elect a President, and twice within a six-year period to vote for two Senators; additionally, various local and municipal elections occur throughout these cycles. In India, the national parliament and state assemblies are legally mandated to have five-year terms; however, as in many parliamentary democracies, the fall of a government can lead to elections being held earlier than the scheduled term.
Some elections are tradionally held together with others, so that a voter on a given poll date might cast votes in multiple elections. For example, the three major American election cycles are synchronized so that every other House election coincides with a Presidential election, and, for each voter, two out of three House elections also involve a senatorial candidate on the ballot. In India, the national and state elections were held simultaneously until 1967 but the premature fall of various governments have caused them to become unsynchronized over the years.
Simultaneous elections reduce cost, time, and effort, and consequently there is a strong economic argument to be made in their favor. The ruling party in India, has proposed a constitutional amendent—“One Nation One Election”—which seeks to mandate the simultaneity of national and state elections. This has been opposed by other parties out of concern that the current Prime Minister’s popularity might increase the state-level dominance of the ruling party. Indeed, the tendency of a popular leader to attract votes for other candidates of the same party is a well-known and much-studied phenomenon in the American context, where it is referred to as a “coattail” or “down-ballot” effect.
A separate development, observed and studied in both the American and Indian contexts, is that of increasing “political polarization” [^1] See, e.g., [7], [21], [40] and the references therein.. This refers to a tendency of voters to identify closely with a political party, with a corresponding reduction in the number of “uncommitted” or “issue-based” voters. At the ballot level this results in more “party-line” votes, with relatively few “split” or “cross-over” votes. While the outcome of an election depends on who shows up to vote, and how they choose to vote; in a polarized society the second factor is much less important. The political parties have a clear idea of who their core supporters are, and much of the election effort is focused on a get-out-the-vote campaign, targeted towards getting these supporters show up on election day.
In this paper we study the likelihood of a “single-party sweep”; namely, given some set of elections, what is the probability that a single party wins all of them? In order for this to be a well-defined question, one must clarify what it means to win an election. In some real-world situations, such as a presidential contest or a two-party parliamentary election, there is an unambiguous notion of what constitutes a “win”; in other settings, there might be several natural candidates.
For example, in the case of Indian state elections, one might be interested in analyzing either the total number of assembly seats won, or the total number of state governments formed by a party. Alternatively one might choose to focus on a single state, examining whether a single party wins majority in that state as well as the national level. The ruling party in India refers to this as “double engine governance”, and portrays it as a favorable outcome that accelerates developmental initiatives within the state [^2] The Indian Railways occasionally operates trains with dual engines to enhance speed and reliability, particularly when handling heavier loads.. For maximal applicability of our analysis, we choose to work with a somewhat abstract notion of a win, subject only to two mild requirements, which fits all these scenarios and many more besides.
Our main finding is that in a country with a polarized electorate, a more simultaneous polling schedule increases the likelihood of a single-party sweep.
We make this statement precise below in three scenarios of increasing generality. Our most general scenario allows for several parties and staggered polling, which is often the case in Indian elections. Although our two motivating examples, the US and India, both involve first-past-the-post voting, this is not assumed in our model, and our results are equally applicable to several other major electoral systems, including party list proportional representation.
Finally, it is worth noting that we actually prove a stronger result below; namely, that probability of a sweep goes up for every party. Thus the effect we describe is systemic in nature and does not depend on the coattails of a popular leader. It stems solely from statistical aggregates of individual voters’ actions, and is best regarded as a macroscopic manifestion of political polarization. Our model brings to light a novel cause for sweeps which is not at odds with other causes discussed in the literature, but rather independent of them.
@@ 1.1 Main results
Country A has two political parties and two national elections, Presidential and Parliamentary; in each election the candidate with the most votes wins, with ties decided by a coin toss. The election commission is considering whether to hold the elections simultaneously or on separate polling dates. Each voter $h$ has a fixed probability $p_{h}$ of voting on any poll date [^3] This assumption is in the spirit of Banzhaf [5] and Penrose [33] (see section 3.1)., and the population is polarized in the sense that each voter prefers the same party for both elections. A sweep is an outcome where one party wins both the presidency and a parliamentary majority.
@@
A simultaneous election makes a sweep more  [^4] Throughout, “more” is to be understood in the weak sense of “greater than or equal to”. likely.
Country B has several political parties, and several elections with different notions of what constitutes a “win”. However the following two conditions hold.
(a) Each election can be “won” by at most one party.
(b) For any turnout in any election, an extra vote for some party cannot decrease its win probability [^5] We speak of “win probabilities” since, for example, tied vote counts might lead to coin tosses., nor can it increase the win probability of a rival party.
The election commission is exploring various polling schedules. We will say that schedule $\pi$ is coarser than schedule $\pi^{\prime}$ if any two elections that are simultaneous (have the same poll date) in $\pi^{\prime}$ are also simultaneous in $\pi$. As before, each voter $h$ has a fixed probability $p_{h}$ of voting on any poll date and always votes for the same party. A sweep is an outcome in which one party wins all the elections.
@@
A coarser polling schedule makes a sweep more likely.
Country C has a setup similar to country B. However the election commission is considering staggered schedules in which each election might be spread out over several days, with different poll dates for different regions. We will say that a staggered schedule $\Pi$ is coarser than $\Pi^{\prime}$ if, for every voter $h$, any two elections that are simultaneous for $h$ in $\Pi^{\prime}$ are also simultaneous for $h$ in $\Pi$.
@@
A coarser staggered schedule makes a sweep more likely.
In a country with many political parties, it often happens that several parties come together in an alliance in order to form a government. Here one can distinguish two kinds of alliances – a pre-poll alliance which involves seat sharing, and a post-poll alliance in which a group of parties comes together after having fought the elections separately, perhaps competing against each other for more seats. In this case, one may ask whether Theorem C continues to hold if we treat each alliance as single party. In section 3.3 we show that is indeed the case for many real-world electoral systems.
Regarding the proofs of these results, it is evident that Theorem C implies Theorem B, which in turn implies Theorem A. We will deduce Theorem C from Theorem D, a mathematical result that we formulate and prove below. Theorem D also yields a new proof of the Harris inequality – a well-known correlation inequality with many mathematical applications [2, 19]. As our proof shows, Theorem D may be regarded as a generalization of the Harris inequality to $n$ functions. [^6] The Harris inequality is a special case of the still more general FKG inequality [17]. One of the authors of the present paper has proposed a different generalization of the FKG inequality to $n$ functions and obtained some partial results in this direction. However the general problem remains open, and seems much harder than the generalization considered here; see [25, 36, 37, 38, 39].
@@ 1.2 Structure of the paper
In section 2 we formulate and prove Theorem D, and deduce Theorem C. We also provide a new proof of the Harris inequality, formulated here as Theorem E.
In Section 3.1 we compare our model of random voter turnout to other models of voter behavior. In Section 3.2 we discuss the applicability of our results to various electoral systems that are in use around the world, and have been studied extensively in the theory of Social Choice [3, 16]. In Section 3.3 we extend Theorem C to party alliances, and in Section 3.4 we contrast our anaysis of sweeps to the earlier literature. Finally, in Sections 3.5 and 3.6 we discuss the implications of our results for elections in the USA and India, respectively.
2 Proofs
We will deduce Theorem C from Theorem D, which we now formulate and prove.
Let $L=\{1,\ldots,n\}$ be a finite set, let $\pi$ be a partition of $L$, and let $\mathcal{L}$ be the power set of $L$. Given $0\leq p\leq 1$, construct a random subset $S\in\mathcal{L}$ as follows: for each $M$ in $\pi$, toss a coin with probability $p$ of heads; if heads then include $M$ in $S$.
@@ 2 Proofs
We write $\mu_{\pi,p}$ for the probability measure on $\mathcal{L}$ such that $\mu_{\pi,p}(S)$ is the probability of obtaining the subset $S$ by the above random procedure.
Let $H=\{1,\ldots,m\}$ be another finite set, and let $\mathcal{H}$ and $\mathcal{T}$ be the power sets of $H$ and $H\times L$. Then we have bijections $\lambda:\mathcal{L}^{m}\rightarrow\mathcal{T},\;\theta:\mathcal{H}^{n}\rightarrow\mathcal{T}$, given by
$\lambda(L_{1},\ldots,L_{m})=\bigcup\nolimits_{h}\{(h,l):l\in L_{h}\},\quad\theta(H_{1},\ldots,H_{n})=\bigcup\nolimits_{l}\{(h,l):h\in H_{l}\}.$
@@
We say that a (real-valued) function $f$ on $\mathcal{H}$ is increasing (resp. decreasing) at $h\in H$, if for all $S\subset H$ one has $f(S\cup\{h\})-f(S)\geq 0$ (resp. $\leq 0$).
We say that $F=(f_{1},\ldots,f_{n})$ is an aligned tuple of functions on $\mathcal{H}$ if, for each $h\in H$, either all the $f_{l}$ are increasing at $h$, or all are decreasing at $h$.
Let $\Pi=(\pi_{1},\ldots,\pi_{m})$, $\mathbf{p}=(p_{1},\ldots,p_{m})$, $F=(f_{1},\ldots,f_{n})$ be tuples such that each $\pi_{h}$ is a partition of $L$, each $p_{h}$ is in $[0,1]$, and each $f_{l}$ is a function on $\mathcal{H}$. We write $\mu_{\Pi,\mathbf{p}}$ and $\bar{F}$ for the product measure and function on $\mathcal{T}$ induced by $\lambda$ and $\theta$:
$\mu_{\Pi,\mathbf{p}}(L_{1},\ldots,L_{m}):=\mu_{\pi_{1},p_{1}}(L_{1})\cdots\mu_{\pi_{m},p_{m}}(L_{m}),\;\bar{F}(H_{1},\ldots,H_{n}):=f_{1}(H_{1})\cdots f_{n}(H_{n}).$
Definition 3
For $\Pi$, $\mathbf{p}$, $F$ as above, we define $E(\Pi,\mathbf{p},F):=\sum\nolimits_{T\in\mathcal{T}}\bar{F}(T)\mu_{\Pi,\mathbf{p}}(T).$
@@
Suppose $F=(f_{1},\ldots,f_{n})$ is an aligned tuple of nonnegative functions on $\mathcal{H}$. If $\;\Pi^{\prime}$ is coarser than $\Pi$ then we have $E(\Pi^{\prime},\mathbf{p},F)\geq E(\Pi,\mathbf{p},F)$ for all $\mathbf{p}$.
Proof. It suffices to prove $E(\Pi^{\prime},\mathbf{p},F)\geq E(\Pi,\mathbf{p},F)$ for the case where
  • 1)
@@
where $p=p_{1}$. Now by an elementary calculation we get
@@
Theorem C follows immediately from Theorem D.
Proof of Theorem C. Let $H=\{1,\ldots,m\}$ be the set of voters and let $L=\{1,\ldots,n\}$ be the set of elections. A voter turnout, $T\subset H\times L,$ is the set of pairs $(h,l)$ such that voter $h$ has voted in election $l$. The staggered schedules $\Pi$ and the voter probabilities $\mathbf{p}=(p_{h})$ induce a probability measure on voter turnouts, which is seen to be precisely $\mu_{\Pi,\mathbf{p}}$. We will prove that if we replace $\Pi$ by a coarser partition then the probability of a sweep goes up for every party.
Fix a party $s$ and define $F=(f_{1},\ldots,f_{n})$ where $f_{i}(H^{\prime})$ is the probability that party $s$ wins election $i$ when $H^{\prime}$ is the set of votes cast. Then the $f_{i}$ are non-negative and aligned by our assumptions. Now the probability of a sweep by party $s$ is precisely $E(\Pi,\mathbf{p},F)$, and so the result follows by Theorem D.
Theorem D also gives a new proof of the Harris correlation inequality [19]. For $\mathbf{p}=(p_{1},\ldots,p_{m})$, define a probabilty measure and expectation on $\mathcal{H}$ as follows:
$\mu(S)=\mu_{\mathbf{p}}(S)=\prod\nolimits_{h\in S}(p_{h})\prod\nolimits_{h\notin S}(1-p_{h}),\quad E(f,\mu)=\sum\nolimits_{S\in\mathcal{H}}f(S)\mu(S).$
@@
If $f_{1},f_{2}$ are increasing functions on $\mathcal{H}$ then we have
$E(f_{1}f_{2},\mu)-E(f_{1},\mu)E(f_{2},\mu)\geq 0.$ (1)
Proof. This is due to Harris [19], but it is also an immediate corollary of the case $n=2$ of Theorem D as we now explain. First, if we add a constant to $f_{1}$ or $f_{2}$ then the left side of (1) is unchanged, thus we may assume without loss of generality that $f_{1}$ and $f_{2}$ are non-negative. Evidently then the pair $F=(f_{1},f_{2})$ is aligned and non-negative. Now let $\Pi=(\pi_{1},\ldots,\pi_{m})$ and $\Pi^{\prime}=(\pi^{\prime}_{1},\ldots,\pi^{\prime}_{m})$ where $\pi_{h}=\{\{1\},\{2\}\}$ and $\pi^{\prime}_{h}=\{\{1,2\}\}$ for all $h$, then we have
$E(\Pi^{\prime},\mathbf{p},F)=E(f_{1}f_{2},\mu),\quad E(\Pi,\mathbf{p},F)=E(f_{1},\mu)E(f_{2},\mu).$
Since $\Pi^{\prime}$ is coarser than $\Pi$, Theorem D implies (1) as desired.
Thus Theorem D can be regarded as an $n$-function generalization of teh Harris inequality.